# Sensor-fusion based Prognostics for Deep-space Habitats Exhibiting Multiple Unlabeled Failure Modes

### v3

@@ 5 Online Diagnosis and Prognostics
In this section, a subsystem(s) in deep-space habitat is considered to be degrading in real-time. Our goal is to predict the time remaining before the unit fails. We begin by showing how to fuse the signals from multiple sensors into a parsimonious set of features. Then, we demonstrate how we utilize the fused sensor information to i) diagnose the active failure mode of the degrading unit, in real-time, and ii) predict its RUL.
@@ 5.1 Sensor Fusion: MFPCA
Without loss of generality, let us redefine the sensor index as $\{1,2,\ldots,\mathcal{P}\}$, where $\mathcal{P}$ is number of sensors we seek to fuse. Now assume that we have monitored the degrading subsystem(s) for a duration given by the time interval $[0,t^{*}]$. Since we are interested in predicting the remaining lifetime, we only use signals from units with failure times greater than $t^{*}$. We define this set as the “$t^{*}-$training” set, whose indices we denote by the set $\mathcal{I}_{t^{*}}$, which has cardinality $N_{t^{*}}$. Furthermore, we smooth the degradation signals using the ’rloess’ method as described in [52]. This method uses weighted least squares with a 2nd-order polynomial for smoothing. The weights come from a bisquare weight function requiring a specified bandwidth parameter. This method also identifies outliers and assigns them weights of zero to improve signal mean estimation. To perform MFPCA, we model the $i$th smoothed multivariate signal evaluated at time $t\in[0,t^{*}]$ as having mean $\boldsymbol{\mu}(t)\in\mathcal{R}^{\mathcal{P}}$ and $\mathcal{P}\times\mathcal{P}$ block covariance function $\mathcal{C}(t,t^{\prime})$, where $\mathcal{C}_{j,j^{\prime}}(t,t^{\prime})=cov(v_{j}(t),v_{j^{\prime}}(t^{\prime}))$. Using Mercer’s theorem [49], we can represent the covariance function as $\mathcal{C}(t,t^{\prime})=\sum_{h=1}^{\infty}\eta_{h}\boldsymbol{\varpi}_{h}(t)\boldsymbol{\varpi}_{h}(t^{\prime})^{T}$, where $(\eta_{h},\boldsymbol{\varpi}_{h}(t))$ are the $h$th eigenvalue-eigenfunction pair and $\eta_{1}\geq\eta_{2}\geq\dots$. Using the first $H$ multivariate eigenfunctions, we can represent the multivariate signal by Eq 13 where the MFPC-scores $\zeta$’s are computed according to Eq 14.
$\boldsymbol{s}_{i}(t)=\boldsymbol{\mu}(t)+\sum_{h=1}^{H}\zeta_{i,h}\boldsymbol{\varpi}_{h}(t)+\boldsymbol{\epsilon}_{i}(t),\vspace*{-1.2em}$ (13)
[…]
@@ 5.2 Failure Mode Diagnosis: K-nearest Neighbors
In order to diagnose the failure mode of a degrading unit at time $t^{*}$, we apply MFPCA on the $t^{*}$ training dataset. Next, $K$ nearest neighbors (KNN) is applied to the standardized MFPC-scores to identify the active failure mode. Here, the distance metric of the KNN classifier is defined using the Euclidean distance between the MFPC-scores. Given this diagnosis, we now develop a model for predicting the unit RUL.
@@ 5.3.1 Time-Varying Functional Regression
When predicting RUL, we are often faced with updating the relationship between time-to-failure and the current values of the extracted signal features (MFPC-scores). The predictor trajectories up to the current time are represented by time-varying MFPC-scores, which are continuously updated as time progresses and are considered to be time-varying predictor variables for the functional regression model. This is referred to as time-varying functional regression. It was first proposed by [53] and has been used for prognostics in [54, 20]. In our problem setting, let $k^{*}$ denote the diagnosed failure mode and $\mathcal{I}_{t^{*}}^{k^{*}}\subset\mathcal{I}_{t^{*}}$ denote the set of indices for training units that failed due to failure mode $k^{*}$, but survived up to time $t^{*}$. Furthermore, let $\mathcal{P}_{k^{*}}\subset\{1,2,...,\mathcal{P}\}$ denote the set of sensors selected (identified as most informative) for failure mode $k^{*}$. To predict RUL, we fit a functional regression model as shown in Eq (15), where $y_{i}=\ln{TTF_{i}}$, $\varphi_{0}^{k^{*}}$ is the bias term, $\boldsymbol{\varphi}^{k^{*}}(t)\in\mathcal{R}^{|\mathcal{P}_{k^{*}}|}$ is the multivariate coefficient function, and $\boldsymbol{s}_{i}^{k^{*}}(t)$ is the multivariate signal consisting only of sensors significant to failure mode $k^{*}$.
[…]
@@
$\boldsymbol{\varphi}(t)=\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{k}^{k^{*}}(t)\vspace*{-0.5em}$ (16)
@@
$\boldsymbol{s}_{i}^{k^{*}}(t)=\boldsymbol{\mu}^{k^{*}}(t)+\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)\vspace*{-0.5em}$ (17)
@@
$\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(y_{i}-c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\vspace*{-1em}$ (18)
@@
$c_{0}^{k^{*}}=\varphi_{0}^{k^{*}}+\int_{0}^{t^{*}}\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)^{T}\boldsymbol{\mu}^{k^{*}}(t)\vspace*{-0.5em}$ (19)
Weighted Version: We note that the offline sensor selection algorithm may generate some outliers, which can impact the accuracy of the regression problem. These outliers become more apparent after identifying the active failure mode and the MFPC-scores to estimate the time-varying functional regression model. To address this issue, we use weighted least squares to estimate the time-varying functional regression model. To derive the weights, we first measure the Euclidean distance between the MFPC score corresponding to a unit $i$ and the centroid of the cluster of the active failure mode. We then use the reciprocal of that distance to define the corresponding weight for the unit, $w_{i}$, of that unit. If the unit’s failure is misclassified, $w_{i}$ would be large, and thus, unit $i$ is assigned a smaller weight in the regression problem. Next, We regress the standardized $\ln{TTF}$ on the MFPC-scores using a penalized weighted linear regression shown in Eq (20). For variable selection, we include the lasso penalty in the optimization problem. To ensure a regression coefficient represents the significance of their respective predictor to the prediction of the failure time, Eq (20) should be read such that $\zeta_{i,h}^{k^{*}}$ is standardized.
$\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(\sqrt{w_{i}}y_{i}-\sqrt{w_{i}}c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\sqrt{w_{i}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}+\lambda\sum_{h=1}^{H_{k^{*}}}\lVert c_{h}^{k^{*}}\rVert$ (20)
@@ 5.3.2 RUL prediction
$\begin{split}{RUL}_{test}=\exp&\bigg{(}\Big{[}c_{0}^{k^{*}}+\sum_{k=1}^{H_{k^{*}}}\zeta_{test,h}^{k^{*}}c_{k}^{k^{*}}\Big{]}\times\\ &\Big{[}(1/(N-1))\sum_{i=1}^{N}\big{(}\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big{)}^{2}\Big{]}\\ &+(1/N)\sum_{i=1}^{N}\ln{TTF_{i}}\bigg{)}-t^{*}\end{split}$ (21)
@@ 6 Case Study 1: Simulated Dataset
We present a simulation study that mimics the operational and diagnostic challenges of a deep-space habitat. In such systems, critical subsystems like life support, power generation, etc., are monitored by numerous sensors, and degradation can occur through multiple unknown failure modes. Due to the scarcity of in-mission failure data and the infeasibility of performing full-scale failure testing in space, simulation-based validation is pursued in this case study. This study enables us to systematically evaluate our methodology’s ability to: (1) identify latent failure modes in unlabeled multivariate sensor data, (2) select informative sensors specific to each mode, and (3) accurately predict remaining useful life (RUL).
@@ 6.1 Simulation Model
In this Eq, $\epsilon_{i,p}^{k}(t)\sim N\left(0,\left(\sigma_{k}^{p}\right)^{2}\right)$ is a white noise process parameterized by $\sigma_{k}^{p}=\mu_{k}/SNR_{p}^{k}$, where $SNR_{p}^{k}$ is the signal-to-noise for sensor $p$ and failure mode $k$. To account for the correlation between the signal trends of the informative sensors and the underlying degradation process, we generate $\theta_{i,p}^{k}$ from the following conditional distribution $\theta_{i,p}^{k}\mid\theta_{i}^{k}\sim N\left(\mu_{k}\left(1-\sqrt{1-\rho_{p}^{k}}\right),(0.1)^{2}\right)$ such that the correlation, $\rho_{p}^{k}$, is uniformly sampled from the interval $[0.80,0.99]$. The process for generating signals for the noninformative sensors is similar to that of the informative sensors, except the interval from which we sample $\rho_{p}^{k}$ is $[0.1,0.6]$. Finally, we randomly assign a sign to $\theta_{i,p}^{k}$ to allow for signals that increase or decrease in response to degradation.
In this simulation, we seek to test the robustness of our methodology to various levels of signal noise. Therefore, we generate three datasets, each with a unique interval from which to sample $SNR_{p}^{k}$. For $p\in\mathcal{J}_{k}$, $SNR_{p}^{k}$ is sampled uniformly from $[2,5]$, $[5,8]$, and $[8,11]$ for datasets 1, 2, and 3, respectively. For $p\in\mathcal{N}_{k}$, $SNR_{p}^{k}$ is sampled uniformly from $[1,3]$ for all three datasets. Figure 3 displays samples of our sensor signals in a 3$\times$4 grid. Each row displays signals from four sensors (1, 3, 5, 19) for a particular signal-to-noise ratio. Sensor 1 is not informative for either failure mode so no discernible trend is present. Sensors 3 and 5 are informative for FM 2 and FM 1, respectively. For these sensors, one signal trends downward due to degradation while the other signal appears to resemble random noise. Finally, for sensor 19, the signal decreases in response to both failure modes.
@@ 6.2 Failure Mode Labeling and Sensor Selection
For all scenarios, the initial failure mode labels were determined randomly. Table II shows that for the lowest and the highest signal-to-noise ratios, the clustering algorithm was capable of achieving at worst 78.8% accuracy for a particular failure mode. The exception is the dataset [5,8], where the algorithm has low accuracy at clustering. This is due to the tuning parameter combination resulting in minimal coefficient shrinkage that enabled the nonsignificant sensors to influence cluster assignments. A slightly larger $\lambda$ value would decrease this error. For all three models, several sensors that are deemed ”noninformative” were selected. This can be attributed to some of these sensors having high values of $\rho_{p}^{k}$ relative to the specified range of [0.1,0.6]. However, simply reporting the sensors selected does not account for how informative they are. Table 4 reports the four sensors with the highest $l_{2}$-norm of the regression coefficients associated with that sensor. Except for the [5,8] dataset, at least three of the four sensors with the largest $l_{2}$-norm for each failure mode are elements of the set of informative sensors. This indicates that the MGR-ASGL model results in the informative sensors having larger regression coefficients (in magnitude) than the noninformative sensors. Given the selected sensors, we want to test the robustness of the online portion of the methodology to predict remaining life.
@@ 7 Case Study 2: NASA Turbofan Engine Dataset
Figure 5: The system diagram used in Case Study 2 adapted from [55]
To further validate our methodology in a realistic operational context, we evaluate its performance on a benchmark dataset that closely mirrors subsystem degradation in deep-space habitats. Specifically, we apply our approach to the NASA’s Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset [55]. C-MAPSS simulates degradation in the components of turbofan engines that share telemetry and fault characteristics with propulsion and environmental control systems of deep space habitats. C-MAPSS involves multivariate sensor monitoring from healthy operation to system failure, with no explicit failure mode labels and only partial degradation histories for test systems. Figure 5 displays a diagram of this turbofan engine.
In this paper, we utilize C-MAPSS-3, which consists of both a training and a test set. Both sets contain 100 observed engine failures caused by either fan degradation or high-pressure compressor degradation. Furthermore, both sets contain data from 21 sensors, which are listed in Table 5. The signals for these sensors are displayed in Figure 6. These sensors monitor the systems from a good-as-new state to engine failure. However, sensors 1, 5, 6, 10, 16, 18, and 19 contain little to no information and are thus removed from the analysis. While the dataset provides the potential causes of failure, it does not provide the failure labels. Furthermore, the test set consists of systems whose degradation is only partially observed along with their corresponding failure times. Therefore, our goal is to use our methodology to predict the remaining life of each testing unit following the final cycle at which their degradation was observed.

### v4

@@ 5 ONLINE DIAGNOSIS AND PROGNOSTICS
In this section, a system within a deep-space habitat is considered to be degrading in real-time. Our goal is to predict the RUL of the system. We begin by showing how to fuse the signals from multiple sensors into a parsimonious set of features. Then, we demonstrate how we utilize the fused sensor information to i) diagnose the active failure mode of the degrading system in real-time, and ii) predict its RUL.
@@ 5.1 Sensor Fusion: MFPCA
Without loss of generality, let us redefine the sensor index as $\{1,2,\ldots,\mathcal{P}\}$, where $\mathcal{P}$ is number of sensors we seek to fuse. Now assume that we have monitored the degrading system for a duration given by the time interval $0,t^{*}]$. Since we are interested in RUL, we only use signals from training systems with failure times greater than $t^{*}$. We define this set as the $t^{*}$-training” set, whose indices we denote by the set $\mathcal{I}_{t^{*}}$, which has cardinality $N_{t^{*}}$. Furthermore, we smooth the degradation signals using the ’rloess’ method as described in [50]. This method uses weighted least squares with a 2nd-order polynomial for smoothing. The weights come from a bisquare weight function requiring a specified bandwidth parameter. This method also identifies outliers and assigns them weights of zero to improve signal mean estimation. To perform MFPCA, we model the $i$th smoothed multivariate signal evaluated at time $t\in[0,t^{*}]$ as having mean $\boldsymbol{\mu}(t)\in\mathcal{R}^{\mathcal{P}}$ and $\mathcal{P}\times\mathcal{P}$ block covariance function $\mathcal{C}(t,t^{\prime})$, where $\mathcal{C}_{j,j^{\prime}}(t,t^{\prime})=cov(v_{j}(t),v_{j^{\prime}}(t^{\prime}))$. Using Mercer’s theorem [47], we can represent the covariance function as $\mathcal{C}(t,t^{\prime})=\sum_{h=1}^{\infty}\eta_{h}\boldsymbol{\varpi}_{h}(t)\boldsymbol{\varpi}_{h}(t^{\prime})^{T}$, where $(\eta_{h},\boldsymbol{\varpi}_{h}(t))$ are the $h$th eigenvalue-eigenfunction pair and $\eta_{1}\geq\eta_{2}\geq\dots$. Using the first $H$ multivariate eigenfunctions, we can represent the multivariate signal by Eq 13 where the MFPC-scores $\zeta$’s are computed according to Eq 14.
$\boldsymbol{s}_{i}(t)=\boldsymbol{\mu}(t)+\sum_{h=1}^{H}\zeta_{i,h}\boldsymbol{\varpi}_{h}(t)+\boldsymbol{\epsilon}_{i}(t),\vskip-11.99998pt$ (13)
$\zeta_{i,h}=\int_{0}^{t^{*}}(\boldsymbol{s}_{i}(t)-\boldsymbol{\mu}(t))^{T}\boldsymbol{\varpi}_{h}(t)dt\vskip-5.0pt$ (14)
@@ 5.2 Failure Mode Diagnosis: K-nearest Neighbors
In order to diagnose the failure mode of a degrading system at time $t^{*}$, we apply MFPCA on the $t^{*}$ training dataset. Next, K nearest neighbors (KNN) is applied to the standardized MFPC-scores to identify the active failure mode. Here, the distance metric of the KNN classifier is defined using the Euclidean distance between the MFPC-scores. Given this diagnosis, we now develop a model for predicting system RUL.
@@ 5.3.1 Time-Varying Functional Regression
When predicting RUL, we are often faced with updating the relationship between time-to-failure and the current values of the extracted signal features (MFPC-scores). The predictor trajectories up to the current time are represented by time-varying MFPC-scores, which are continuously updated as time progresses and are considered to be time-varying predictor variables for the functional regression model. This is referred to as time-varying functional regression. It was first proposed by [51] and has been used for prognostics in [52, 18]. In our problem setting, let $k^{*}$ denote the diagnosed failure mode and $\mathcal{I}_{t^{*}}^{k^{*}}\subset\mathcal{I}_{t^{*}}$ denote the set of indices for training systems that failed due to failure mode $k^{*}$, but survived up to time $t^{*}$. Furthermore, let $\mathcal{P}_{k^{*}}\subset\{1,2,...,\mathcal{P}\}$ denote the set of sensors selected (identified as most informative) for failure mode $k^{*}$. To predict RUL, we fit a functional regression model as shown in Eq (15), where $y_{i}=\ln{TTF_{i}}$, $\varphi_{0}^{k^{*}}$ is the bias term, $\boldsymbol{\varphi}^{k^{*}}(t)\in\mathcal{R}^{|\mathcal{P}_{k^{*}}|}$ is the multivariate coefficient function, and $\boldsymbol{s}_{i}^{k^{*}}(t)$ is the multivariate signal consisting only of sensors significant to failure mode $k^{*}$.
$\min_{\varphi_{0}^{k^{*}},\boldsymbol{\varphi}^{k^{*}}(t)}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(y_{i}-\varphi_{0}-\int_{0}^{t^{*}}\boldsymbol{\varphi}^{k^{*}}(t)^{T}\boldsymbol{s}_{i}^{k^{*}}(t)dt\right)^{2}\vskip-4.62497pt$ (15)
@@
$\boldsymbol{\varphi}(t)=\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{k}^{k^{*}}(t)\vskip-5.0pt$ (16)
@@
$\boldsymbol{s}_{i}^{k^{*}}(t)=\boldsymbol{\mu}^{k^{*}}(t)+\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)\vskip-5.0pt$ (17)
@@
$\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(y_{i}-c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\vskip-10.00002pt$ (18)
@@
$c_{0}^{k^{*}}=\varphi_{0}^{k^{*}}+\int_{0}^{t^{*}}\sum_{h=1}^{H_{k^{*}}}c_{h}^{k^{*}}\boldsymbol{\varpi}_{h}^{k^{*}}(t)^{T}\boldsymbol{\mu}^{k^{*}}(t)\vskip-5.0pt$ (19)
Weighted Version: The offline sensor selection algorithm is unlikely to cluster failures perfectly. Therefore, when fitting the regression model, it is possible that sample systems that failed due to other failure modes are included. These systems may present themselves as outliers, compromising RUL prediction accuracy. To address this issue, we use weighted least squares to fit the time-varying functional regression model. To derive the weights, we first measure the Euclidean distance between the MFPC score of each training system to the centroid of the cluster of the diagnosed failure mode. Then we calculate the reciprocal of that distance to define the corresponding weights for each system in the training set. If a training system’s failure mode is misclassified, its corresponding weight $w_{i}$ would be large, and thus, training system $i$ would be assigned a smaller weight in the regression problem. Next, we regress the standardized $\ln{TTF}$ on the MFPC-scores using a penalized weighted linear regression shown in Eq (20). For variable selection, we include the lasso penalty in the optimization problem. To ensure a regression coefficient represents the significance of their respective predictor to the prediction of the failure time, Eq (20) should be read such that $\zeta_{i,h}^{k^{*}}$ is standardized.
$\begin{split}\min_{c_{0}^{k^{*}},c_{1}^{k^{*}},\ldots,c_{H_{k^{*}}}^{k^{*}}}&\sum_{i\in\mathcal{I}_{t^{*}}^{k^{*}}}\left(\sqrt{w_{i}}y_{i}-\sqrt{w_{i}}c_{0}^{k^{*}}-\sum_{h=1}^{H_{k^{*}}}\sqrt{w_{i}}\zeta_{i,h}^{k^{*}}c_{h}^{k^{*}}\right)^{2}\\ &+\lambda\sum_{h=1}^{H_{k^{*}}}\lVert c_{h}^{k^{*}}\rVert\end{split}$ (20)
@@ 5.3.2 RUL prediction
$\begin{split}{RUL}_{test}=\exp&\bigg(\Big[c_{0}^{k^{*}}+\sum_{k=1}^{H_{k^{*}}}\zeta_{test,h}^{k^{*}}c_{k}^{k^{*}}\Big]\times\Big[(1/(N-1))\\ &\sum_{i=1}^{N}\big(\ln{TTF}_{i}-(1/N)\sum_{i=1}^{N}\ln{TTF}_{i}\big)^{2}\Big]\\ &+(1/N)\sum_{i=1}^{N}\ln{TTF_{i}}\bigg)-t^{*}\end{split}$ (21)
@@ 6 CASE STUDY 1: SIMULATED DATASET
We present a simulation study that mimics the operational and diagnostic challenges of a deep-space habitat. In such systems, critical subsystems like life support, power generation, etc., are monitored by numerous sensors, and degradation can occur through multiple unknown failure modes. Due to the scarcity of in-mission failure data and the infeasibility of performing full-scale failure testing in space, simulation-based validation is pursued in this case study. This study enables us to systematically evaluate our methodology’s ability to: (1) identify latent failure modes in unlabeled multivariate sensor data, (2) select informative sensors specific to each mode, and (3) accurately predict RUL.
@@ 6.1 Simulation Model
In this equation, $\epsilon_{i,p}^{k}(t)\sim N\left(0,\left(\sigma_{k}^{p}\right)^{2}\right)$ is a white noise process parameterized by $\sigma_{k}^{p}=\mu_{k}/SNR_{p}^{k}$, where $SNR_{p}^{k}$ is the SNR for sensor $p$ and failure mode $k$. To account for the correlation between the signal trends of the informative sensors and the underlying degradation process, we generate $\theta_{i,p}^{k}$ from the following conditional distribution $\theta_{i,p}^{k}\mid\theta_{i}^{k}\sim N\left(\mu_{k}\left(1-\sqrt{1-\rho_{p}^{k}}\right),(0.1)^{2}\right)$ such that the correlation, $\rho_{p}^{k}$, is uniformly sampled from the interval $[0.80,0.99]$. The process for generating signals for the noninformative sensors is similar to that of the informative sensors, except the interval from which we sample $\rho_{p}^{k}$ is $[0.1,0.6]$. Finally, we randomly assign a sign to $\theta_{i,p}^{k}$ to allow for signals that increase or decrease in response to degradation.
[…]
@@ 6.2 Failure Mode Labeling and Sensor Selection
For all scenarios, the initial failure mode labels were determined randomly. Table II shows that for the lowest and the highest SNRs, the clustering algorithm was capable of achieving at worst 78.8% accuracy for a particular failure mode. The exception is the dataset [5,8], where the algorithm has low accuracy at clustering. This is due to the tuning parameter combination resulting in minimal coefficient shrinkage that enabled the nonsignificant sensors to influence cluster assignments. A slightly larger $\lambda$ value would decrease this error. For all three models, several sensors that are deemed ”noninformative” were selected. This can be attributed to some of these sensors having high values of $\rho_{p}^{k}$ relative to the specified range of [0.1,0.6]. However, simply reporting the sensors selected does not account for how informative they are. Table 4 reports the four sensors with the highest $l_{2}$-norm of the regression coefficients associated with that sensor. Except for the [5,8] dataset, at least three of the four sensors with the largest $l_{2}$-norm for each failure mode are elements of the set of informative sensors. This indicates that the MGR-ASGL model results in the informative sensors having larger regression coefficients (in magnitude) than the noninformative sensors. Given the selected sensors, we want to test the robustness of the online portion of the methodology to predict remaining life.
6.3 RUL Prediction
 $\text{ Relative Error }=\frac{\mid\text{Estimated Life}-\text{Actual Life}\mid}{\text{Actual Life}}\times 100\%$ (24)
Figure 4: Prediction error (in $\%$) for all test systems of case study 1 under three different SNRs i.e., [2, 5], [5, 8], and [8, 11]
Figure 5: The system diagram used in Case Study 2 adapted from [53]
Figure 7: Average Prediction error ($\%$) in case study 2 with comparison to [38], and [37]
Figure 8: Prediction error ($\%$) for different RUL intervals in case study 2 with comparison to [37]
@@ 7 CASE STUDY 2: NASA TURBOFAN ENGINE DATASET
To further validate our methodology in a realistic operational context, we evaluate its performance on a benchmark dataset that closely mirrors subsystem degradation in deep-space habitats. Specifically, we apply our approach to NASA’s Commercial Modular Aero-Propulsion System Simulation (C-MAPSS) dataset [53]. C-MAPSS simulates degradation in the components of turbofan engines that share telemetry and fault characteristics with propulsion and environmental control systems of deep space habitats. C-MAPSS involves multivariate sensor monitoring from healthy operation to system failure, with no explicit failure mode labels and only partial degradation histories for test systems. Figure 5 displays a diagram of this turbofan engine.
In this paper, we utilize C-MAPSS-3, which consists of both a training and a test set. Both sets contain 100 observed engine failures caused by either fan degradation or high-pressure compressor degradation. Furthermore, both sets contain data from 21 sensors, which are listed in Table 5. The signals for these sensors are displayed in Figure 6. These sensors monitor the systems from a good-as-new state to engine failure. However, sensors 1, 5, 6, 10, 16, 18, and 19 contain little to no information and are thus removed from the analysis. While the dataset provides the potential causes of failure, it does not provide the failure labels. Furthermore, the test set consists of systems for which degradation is only partially observed, together with their corresponding failure times. Therefore, our goal is to use our methodology to predict the remaining life of each testing unit following the final cycle at which their degradation was observed.
@@
7.1 Comparison to Baselines
To analyze our methodology, we compare its performance to that of [37] Chehade et al. (2018), whose methodology assumes knowledge of the failure modes but does not perform sensor selection (outside of removing the blank sensors), and [38] Li et al. (2023), whose methodology assumes knowledge of a fraction of the training set and does perform sensor selection. Figure 7 displays the relative prediction error averaged over all test systems with RUL at most 100, 80, 60, 40, and 20 cycles. From this figure, our methodology becomes more accurate as the RUL decreases. Furthermore, our methodology is more accurate than both when making longer predictions. For test units whose RUL is at most 20, our accuracy is between that of [37] and [38]. Only at RUL at most 40, does our methodology perform worse than both. To further analyze our methodology, we refer to Figure 8, which displays boxplots showing the distribution of the relative prediction error for test units with remaining life in the following intervals: [0,25],[26,50],[51,75],[76,100],[101,125],[126,150][0,25],[26,50],[51,75],[76,100],[101,125],[126,150]. The results indicate that our methodology achieves a lower median relative error than [37] in all intervals while maintaining comparable variability. However, the variability for systems with RUL between 25 and 50 is higher for our methodology than for [37]. This indicates a targeted area for improvement of our methodology. However, the general trend is that sensor selection resulted in a more accurate RUL prediction than no sensor selection. Furthermore, we were capable of achieving more accurate RUL prediction than both baselines despite not having information about the actual failure mode.